What is the amount of six equal annual deposits that can provide ten annual withdrawals, where a first withdrawal of $17223 is made at the end of year seven and subsequent withdrawals increase at $846 over the previous year, at an interest rate of 5% compounded annually? Enter you answer as follows: 12345 Round your answer. Do not use a dollar sign ($), any commas() or a decimal point ()
Answer #1

1. Finding PV of All Withdrawals

7 17223      12,240.06
8 18069      12,229.81
9 18915      12,192.78
10 19761      12,131.54
11 20607      12,048.49
12 21453      11,945.83
13 22299      11,825.64
14 23145      11,689.80
15 23991      11,540.08
16 24837      11,378.12
PV of All Withdrawals 1,19,222.14

Next: PV of All Deposits= x/(1.05)*1 + x/(1.05)^2 + .... and so on

This yields PV of all deposits= x(1-0.7452)/(1-0.9523) = 5.3417x
Now, 5.3417x= 119222.14

which yields x= 22319.14